jaoun:

LC)Simplify negative 5 and 1 over 6 − negative 8 and 1 over 2. Which of the following is correct? negative 13 and 2 over 3 negative 3 and 1 over 3 3 and 1 over 3 13 and 2 over 3

jhonyy9:

\[-5\frac{ 1 }{ 6 } - 8\frac{ 1 }{ 2 } = ?\] do you think on this ?

Vocaloid:

old question, will answer so this can be closed rewrite as: -5 - (1/6) - 8 - (1/2) subtracting the whole numbers only: - 5 - 8 = -13 subtracting the fractions only:- (1/6) - (1/2) = -(1/6 + 1/2) = -(1/6 + 3/6) = -(4/6) = -2/3 combining: -13 and 2/3